OpenStudy (anonymous):

the grocery store sells $4.00 a pound and pomegrants for $2.75 a pound. Choose an equation in standard form for the weights of data and pomegrants p that a customer could buy with 12 dollars . A. 4p+ 2.75d = 12 B. 4d= 2.75p +12 C. 4d + 2.75 = 12 D. 4+ 2.75 = d

OpenStudy (anonymous):

the cost of both data and pomegrants together must equal $12

OpenStudy (anonymous):

and you multiply data with the data price, pomgreant with pomgreant price to get pounds of fruit -> price

OpenStudy (anonymous):

"pomegrants for $2.75 a pound" so if you have one pound - p=1, you will have to pay 1 x $2.75 or p x $2.75

OpenStudy (anonymous):

why would you multiply? and how did you get that

OpenStudy (anonymous):

you multiply because it's a price per pound \[pomegrant:\frac{ $2.75 }{ pound }\] if you want the price for some weight, you multiply this per pound price with the actual pound, ex. you buy 2 pounds: \[\frac{ $2.75 }{ pound } \times 2 pound=$5.5\]

OpenStudy (anonymous):

wait so what would the answer be for this ?

OpenStudy (anonymous):

did you copy a-D precisely ?

OpenStudy (anonymous):

i copy d as it was 4 the answers

OpenStudy (anonymous):

you multiply because it is a price for one pound so if you have two pounds, you pay two times that one pound price if you only buy half a pound, you pay half the one pound price

OpenStudy (anonymous):

because I would expect the eq to be 12 = 4 d + 2.75 p ....

OpenStudy (anonymous):

wait im lost here

OpenStudy (anonymous):

ok tell me whats unclear

OpenStudy (anonymous):

so would it be C

OpenStudy (anonymous):

yes but the C does have only 2.75 and no p

OpenStudy (anonymous):

yes does i forgot to put it in there

OpenStudy (anonymous):

ok , then that's the answer :)

OpenStudy (anonymous):

thank you